# Cascading Multi-Agent Anomaly Detection in Surveillance Systems via Vision-Language Models and Embedding-Based Classification

### v1

@@ 1 Introduction
Nam dui ligula, fringilla a, euismod sodales, sollicitudin vel, wisi. Morbi auctor lorem non justo. Nam lacus libero, pretium at, lobortis vitae, ultricies et, tellus. Donec aliquet, tortor sed accumsan bibendum, erat ligula aliquet magna, vitae ornare odio metus a mi. Morbi ac orci et nisl hendrerit mollis. Suspendisse ut massa. Cras nec ante. Pellentesque a nulla. Cum sociis natoque penatibus et magnis dis parturient montes, nascetur ridiculus mus. Aliquam tincidunt urna. Nulla ullamcorper vestibulum turpis. Pellentesque cursus luctus mauris. Nulla malesuada porttitor diam. Donec felis erat, congue non, volutpat at, tincidunt tristique, libero. Vivamus viverra fermentum felis. Donec nonummy pellentesque ante. Phasellus adipiscing semper elit. Proin fermentum massa ac quam. Sed diam turpis, molestie vitae, placerat a, molestie nec, leo. Maecenas lacinia. Nam ipsum ligula, eleifend at, accumsan nec, suscipit a, ipsum. Morbi blandit ligula feugiat magna. Nunc eleifend consequat lorem. Sed lacinia nulla vitae enim. Pellentesque tincidunt purus vel magna. Integer non enim. Praesent euismod nunc eu purus. Donec bibendum quam in tellus. Nullam cursus pulvinar lectus. Donec et mi. Nam vulputate metus eu enim. Vestibulum pellentesque felis eu massa.
2 Introduction
@@
Anomaly detection in surveillance video has become an indispensable component of intelligent monitoring systems, as large-scale camera deployments generate continuous streams that cannot be reviewed manually. Effective systems must satisfy two conflicting requirements: real-time responsiveness, to react to critical events with minimal latency, and semantic interpretability, to ensure that alerts correspond to meaningful anomalies [6]. Conventional methods achieve only partial success. Object-detection pipelines, including recent high-speed variants of the YOLO family, provide rapid recognition but are limited to predefined categories and fail to generalize to complex or unseen events  [15]. Reconstruction-based methods, such as convolutional autoencoders, detect deviations from learned standard patterns without requiring explicit labels, but are often unstable under illumination changes, camera noise, or dynamic backgrounds  [7].
@@
Recent surveys emphasize the potential of vision-language models (VLMs) for video anomaly detection, highlighting their ability to generate natural-language descriptions and capture contextual semantics  [12]. Emerging frameworks, such as LLaVA-NeXT-Interleave, extend these capabilities to multi-image and video scenarios, while weakly supervised methods, like VadCLIP, demonstrate how textual prototypes can align with video features to detect anomalies without exhaustive annotations  [17]. Further advances in unsupervised VLM-assisted anomaly detection explore promising applications in open-world surveillance contexts. Nevertheless, the high inference latency and computational footprint of VLMs constrain their direct use in real-time deployments. The literature increasingly points toward cascading architectures as a viable solution  [18]. Lightweight detectors handle routine inputs at the edge, while only ambiguous or complex cases escalate to computationally intensive reasoning modules. This mirrors trends in networking and multi-agent surveillance, where event-driven responsiveness is complemented by background monitoring and infrastructure health checks. Such designs not only improve efficiency but also align with requirements for privacy preservation and scalability in distributed multi-camera systems  [5].
@@
This work introduces a cascading multi-agent architecture that integrates three complementary paradigms: fast object-level detection using YOLOv8, reconstruction-based anomaly scoring using a convolutional autoencoder, and semantic reasoning using a VLM with an embedding-based classifier  [10]. The design employs an event-driven agent that responds to silent alarms and a cyclical monitoring agent that systematically verifies camera streams and configurations. Together, these agents achieve both rapid responsiveness and long-term robustness. The cascading pipeline ensures early exit for routine cases while escalating ambiguous frames for deeper semantic analysis, thus balancing efficiency with interpretability  [8].
@@
Together, these contributions establish a reproducible and generalizable framework that advances anomaly detection in surveillance by unifying multi-agent coordination, cascading efficiency, and semantic reasoning.
@@ 3 Related Work
This review examines recent studies pertinent to surveillance video anomaly detection through three strands: efficient detector-based methods, weakly supervised and vision-language assisted anomaly detection, and domain-agnostic frameworks. The literature search was conducted across Google Scholar, arXiv, IEEE Xplore, and ACM DL using keywords such as video anomaly detection, weak supervision, vision-language models, edge detection vs reconstruction, and multi-camera surveillance. After removing duplicates and inaccessible works, eight papers from 2024-2025 were selected based on relevance, methodological innovation, and performance metrics  [1, 13].
@@
Wu et al. [16] introduced VadCLIP, which adapts CLIP for weakly supervised anomaly detection by aligning video embeddings with textual prototypes. Their methodology reduced reliance on frame-level labels and achieved strong performance on UCF-Crime and XD-Violence. Results confirmed that multimodal alignment improves generalization to unseen events. However, inference latency limited real-time deployment, and the system struggled with fine-grained contextual anomalies. Extending this work, Liu et al. [9] developed VadCLIP++, adding dynamic temporal modules to improve context sensitivity. This resulted in higher detection accuracy and more stable performance under temporal noise. Nevertheless, its computational overhead remained prohibitive for edge-level surveillance, suggesting the need for hybrid pipelines where VLMs are applied selectively.
@@
Compact and efficient architectures have also been proposed. Zhu et al. [20] designed ProDisc-VAD, which introduced a prototype interaction layer and discriminative enhancement loss, achieving AUCs above 97% on ShanghaiTech and 87% on UCF-Crime with only 0.4M parameters. The model demonstrated that weakly supervised anomaly detection can be both accurate and lightweight. Still, its reliance on video-level labels restricted anomaly localization, and semantic reasoning was absent. Gu et al. [3] proposed UniVAD, a training-free few-shot framework that clusters patches and builds graph-based reasoning for cross-domain anomaly detection. Results showed strong generalization across industrial, medical, and surveillance datasets. Yet, UniVAD operated at image level rather than on continuous streams, limiting its applicability for real-time monitoring.
@@
Explainability has been another critical focus. Huang et al. [4] introduced Ex-VAD, which coupled vision-language models with natural-language generation, enabling fine-grained anomaly classification with textual explanations. Empirical evaluations highlighted improvements in interpretability and user trust, though at the cost of increased latency. Ye et al. [19] presented VERA, which verbalizes anomaly reasoning directly from VLMs, offering transparent decision pathways. While valuable for accountability, the method depended on computationally heavy VLM inference and was difficult to scale.
@@
Dataset contributions and surveys have also influenced the field. Zhu et al. [21] released MSAD, a dataset with diverse illumination and scene variations, while also reviewing recent anomaly detection models. Their work emphasized dataset bias as a persistent barrier to real-world robustness. Gao et al. [2] provided a unifying perspective on the evolution of video anomaly detection from deep neural networks to multimodal large models, underscoring the importance of hybrid systems that combine efficiency with contextual reasoning. Sträter et al. [14] introduced GeneralAD, a cross-domain model that simulates distortions and attends to patch features for anomaly localization. Although results were strong across domains, the model required significant computational resources and lacked explicit semantic reasoning.
@@
Wu et al. [16] VadCLIP: CLIP-based VLM Surveillance High latency Basis for selective VLM use
Liu et al. [9] VadCLIP++: temporal modules Surveillance Heavy compute Motivates hybrid cascade
Zhu et al. [20] ProDisc-VAD: prototype + discrim. loss Surveillance Video-level labels Lightweight baseline
Gu et al. [3] UniVAD: few-shot graph reasoning Cross-domain Image-level only Shows generalization limits
Huang et al. [4] Ex-VAD: explainable VLM Surveillance Latency; misclass. Supports interpretable VLM
Ye et al. [19] VERA: verbalized reasoning Surveillance Heavy inference Adds semantic transparency
Zhu et al. [21] MSAD: dataset + review Surveillance Dataset bias Benchmark reference
Gao et al. [2] Survey: DNN to multimodal Broad No pipeline Confirms hybrid trend
Sträter et al. [14] GeneralAD: transformer model Cross-domain High compute; no semantics Illustrates trade-offs
@@ 6.1 Dataset
The evaluation is performed on the UCF-Crime dataset [11], one of the largest benchmarks for real-world video anomaly detection. The dataset comprises 1,900 untrimmed surveillance videos with a total duration of approximately 128 hours. Each video is recorded at 30 frames per second and a resolution of $1280\times 720$, resulting in nearly 13.8 million frames. The average video length is 242 seconds (about 7,260 frames per sequence). The dataset provides 1,610 training videos with video-level labels (normal versus anomalous) and 290 test videos with frame-level temporal annotations. Anomalous intervals are defined by explicit start–end frame indices, and multiple anomalous segments may be present within a single video. Thirteen categories of anomalous events are included: Abuse, Arrest, Arson, Assault, Burglary, Explosion, Fighting, Road Accident, Robbery, Shooting, Stealing, Shoplifting, and Vandalism, along with a large set of normal videos. The dataset is inherently imbalanced, with normal sequences forming the majority and several anomalous classes under-represented (e.g., Shooting, Road Accident). This imbalance necessitates evaluation with macro-averaged metrics such as AUROC and F1-score rather than raw accuracy alone.
@@ Acknowledgments
This work was funded under the PNRR Next Generation EU, pursuant to Article 8 of Ministerial Decree 630/2024, and carried out in collaboration with SPEE Srl, L’Aquila, under the supervision of Professor Giovanni De Gasperis. The research was also supported by the Research and Innovation Project “IMAHS – Intelligent Multi-agent Architecture for Himalaya Supervisor”, funded by Public Intervention 1.1.1.1 of the Abruzzo Regional Programme ERDF 2021–2027, which supports research and innovation projects of companies aligned with the technological domains of the regional Smart Specialization Strategy (RIS3 Abruzzo 2021–2027).

### v2

@@ 1 Introduction
Anomaly detection in surveillance video has become an indispensable component of intelligent monitoring systems, as large-scale camera deployments generate continuous streams that cannot be reviewed manually. Effective systems must satisfy two conflicting requirements: real-time responsiveness, to react to critical events with minimal latency, and semantic interpretability, to ensure that alerts correspond to meaningful anomalies li2024llava . Conventional methods achieve only partial success. Object-detection pipelines, including recent high-speed variants of the YOLO family, provide rapid recognition but are limited to predefined categories and fail to generalize to complex or unseen events  wang2024yolov9 . Reconstruction-based methods, such as convolutional autoencoders, detect deviations from learned standard patterns without requiring explicit labels, but are often unstable under illumination changes, camera noise, or dynamic backgrounds  liu2025networking .
@@
Recent surveys emphasize the potential of vision-language models (VLMs) for video anomaly detection, highlighting their ability to generate natural-language descriptions and capture contextual semantics  neloy2024comprehensive . Emerging frameworks, such as LLaVA-NeXT-Interleave, extend these capabilities to multi-image and video scenarios, while weakly supervised methods, like VadCLIP, demonstrate how textual prototypes can align with video features to detect anomalies without exhaustive annotations  wu2024deep . Further advances in unsupervised VLM-assisted anomaly detection explore promising applications in open-world surveillance contexts. Nevertheless, the high inference latency and computational footprint of VLMs constrain their direct use in real-time deployments. The literature increasingly points toward cascading architectures as a viable solution  wu2024vadclip . Lightweight detectors handle routine inputs at the edge, while only ambiguous or complex cases escalate to computationally intensive reasoning modules. This mirrors trends in networking and multi-agent surveillance, where event-driven responsiveness is complemented by background monitoring and infrastructure health checks. Such designs not only improve efficiency but also align with requirements for privacy preservation and scalability in distributed multi-camera systems  jiang2024vision .
@@
This work introduces a cascading multi-agent architecture that integrates three complementary paradigms: fast object-level detection using YOLOv8, reconstruction-based anomaly scoring using a convolutional autoencoder, and semantic reasoning using a VLM with an embedding-based classifier  liu2025privacy . The design employs an event-driven agent that responds to silent alarms and a cyclical monitoring agent that systematically verifies camera streams and configurations. Together, these agents achieve both rapid responsiveness and long-term robustness. The cascading pipeline ensures early exit for routine cases while escalating ambiguous frames for deeper semantic analysis, thus balancing efficiency with interpretability  liu2025vadclip++ .
@@
Together, these contributions establish a reproducible and generalizable framework that advances anomaly detection in surveillance by unifying multi-agent coordination, cascading efficiency, and semantic reasoning. This work is developed as part of the research project “IMAHS – Intelligent Multi-Agent Architecture for Himalaya Supervisor”, which aims to design scalable multi-agent solutions for intelligent supervision and monitoring. In this context, our proposed cascading multi-agent anomaly detection framework contributes a practical pipeline for real-time surveillance analysis by combining Vision–Language Models with embedding-based classification and multi-stage escalation.
@@ 2 Related Work
This review examines recent studies pertinent to surveillance video anomaly detection through three strands: efficient detector-based methods, weakly supervised and vision-language assisted anomaly detection, and domain-agnostic frameworks. The literature search was conducted across Google Scholar, arXiv, IEEE Xplore, and ACM DL using keywords such as video anomaly detection, weak supervision, vision-language models, edge detection vs reconstruction, and multi-camera surveillance. After removing duplicates and inaccessible works, eight papers from 2024-2025 were selected based on relevance, methodological innovation, and performance metrics  ebidor2024literature ; shelby2008understanding .
@@
Wu et al. wu2024 introduced VadCLIP, which adapts CLIP for weakly supervised anomaly detection by aligning video embeddings with textual prototypes. Their methodology reduced reliance on frame-level labels and achieved strong performance on UCF-Crime and XD-Violence. Results confirmed that multimodal alignment improves generalization to unseen events. However, inference latency limited real-time deployment, and the system struggled with fine-grained contextual anomalies. Extending this work, Liu et al. liu2025 developed VadCLIP++, adding dynamic temporal modules to improve context sensitivity. This resulted in higher detection accuracy and more stable performance under temporal noise. Nevertheless, its computational overhead remained prohibitive for edge-level surveillance, suggesting the need for hybrid pipelines where VLMs are applied selectively.
@@
Compact and efficient architectures have also been proposed. Zhu et al. zhu2025 designed ProDisc-VAD, which introduced a prototype interaction layer and discriminative enhancement loss, achieving AUCs above 97% on ShanghaiTech and 87% on UCF-Crime with only 0.4M parameters. The model demonstrated that weakly supervised anomaly detection can be both accurate and lightweight. Still, its reliance on video-level labels restricted anomaly localization, and semantic reasoning was absent. Gu et al. gu2024 proposed UniVAD, a training-free few-shot framework that clusters patches and builds graph-based reasoning for cross-domain anomaly detection. Results showed strong generalization across industrial, medical, and surveillance datasets. Yet, UniVAD operated at image level rather than on continuous streams, limiting its applicability for real-time monitoring.
@@
Explainability has been another critical focus. Huang et al. huang2024 introduced Ex-VAD, which coupled vision-language models with natural-language generation, enabling fine-grained anomaly classification with textual explanations. Empirical evaluations highlighted improvements in interpretability and user trust, though at the cost of increased latency. Ye et al. ye2025 presented VERA, which verbalizes anomaly reasoning directly from VLMs, offering transparent decision pathways. While valuable for accountability, the method depended on computationally heavy VLM inference and was difficult to scale.
@@
Dataset contributions and surveys have also influenced the field. Zhu et al. zhu2024 released MSAD, a dataset with diverse illumination and scene variations, while also reviewing recent anomaly detection models. Their work emphasized dataset bias as a persistent barrier to real-world robustness. Gao et al. gao2025 provided a unifying perspective on the evolution of video anomaly detection from deep neural networks to multimodal large models, underscoring the importance of hybrid systems that combine efficiency with contextual reasoning. Sträter et al. strater2024 introduced GeneralAD, a cross-domain model that simulates distortions and attends to patch features for anomaly localization. Although results were strong across domains, the model required significant computational resources and lacked explicit semantic reasoning.
@@
Wu et al. wu2024 VadCLIP: CLIP-based VLM Surveillance High latency Basis for selective VLM use
Liu et al. liu2025 VadCLIP++: temporal modules Surveillance Heavy compute Motivates hybrid cascade
Zhu et al. zhu2025 ProDisc-VAD: prototype + discrim. loss Surveillance Video-level labels Lightweight baseline
Gu et al. gu2024 UniVAD: few-shot graph reasoning Cross-domain Image-level only Shows generalization limits
Huang et al. huang2024 Ex-VAD: explainable VLM Surveillance Latency; misclass. Supports interpretable VLM
Ye et al. ye2025 VERA: verbalized reasoning Surveillance Heavy inference Adds semantic transparency
Zhu et al. zhu2024 MSAD: dataset + review Surveillance Dataset bias Benchmark reference
Gao et al. gao2025 Survey: DNN to multimodal Broad No pipeline Confirms hybrid trend
Sträter et al. strater2024 GeneralAD: transformer model Cross-domain High compute; no semantics Illustrates trade-offs
@@ 5.1 Dataset
The evaluation is performed on the UCF-Crime dataset kaggle_ucfcrime , one of the largest benchmarks for real-world video anomaly detection. The dataset comprises 1,900 untrimmed surveillance videos with a total duration of approximately 128 hours. Each video is recorded at 30 frames per second and a resolution of $1280\times 720$, resulting in nearly 13.8 million frames. The average video length is 242 seconds (about 7,260 frames per sequence). The dataset provides 1,610 training videos with video-level labels (normal versus anomalous) and 290 test videos with frame-level temporal annotations. Anomalous intervals are defined by explicit start–end frame indices, and multiple anomalous segments may be present within a single video. Thirteen categories of anomalous events are included: Abuse, Arrest, Arson, Assault, Burglary, Explosion, Fighting, Road Accident, Robbery, Shooting, Stealing, Shoplifting, and Vandalism, along with a large set of normal videos. The dataset is inherently imbalanced, with normal sequences forming the majority and several anomalous classes under-represented (e.g., Shooting, Road Accident). This imbalance necessitates evaluation with macro-averaged metrics such as AUROC and F1-score rather than raw accuracy alone.
@@ Acknowledgments
This work was partially funded under the PNRR Next Generation EU (Article 8, Ministerial Decree 630/2024) and supported by the Research and Innovation Project “IMAHS – Intelligent Multi-agent Architecture for Himalaya Supervisor”, funded through Public Intervention 1.1.1.1 of the Abruzzo Regional Programme ERDF 2021–2027 (RIS3 Abruzzo 2021–2027). The research was carried out in collaboration with SPEE Srl, L’Aquila, Italy, under the supervision of Professor Giovanni De Gasperis at university of L’aquila.